FAR-INFRARED LASER MAGNETIC RESONANCE SPECTROSCOPIC STUDY OF THE ν_{2} BENDING FUNDAMENTAL OF THE CCN RADICAL IN ITS X̃^{2}Π_{r} STATE

\underline{M. D. ALLEN}, K. M. EVENSON, Time and Frequency Division, National Institute of Standards and Technology, Boulder, CO 80303-3328; and J. M. BROWN, The Physical Chemistry Laboratory, Oxford University, South Parks Road, Oxford OX1 3QZ, United Kingdom.

Vibration-rotation transitions between the (010) μ^{2}Σ^{−} - (000) X̃^{2}Π_{r} vibronic states, were recorded using far-infrared laser magnetic resonance (FIR-LMR) spectroscopic techniques. These transitions occur near 200 cm^{−1}for the (010) μ^{2}Σ^{−} - (000) X̃^{2}Π_{1/2} transition and 160 cm^{−1}for the (010) μ^{2}Σ^{−} - (000) X̃^{2}Π_{3/2} transition. This is the first direct measurement of the ν_{2} band of CCN and in conjunction with a fit of optical data, similar to the one found in the paper by Kohguchi et al.,^{a} has resulted in an accurate determination of the bending vibration frequency and the Renner parameter. The data were fit using an N^{2} effective Hamiltonian modified to include the Renner-Teller effect.

aH. Kohguchi, Y. Ohshima, Y. Endo, J. Chem. Phys. 106, 5429 (1997).